OpenStudy (fools101):

I have one more! Logarithmc Equations log (x) + log (x + 9) = 1

OpenStudy (fools101):

please help!

OpenStudy (anonymous):

x=1

OpenStudy (fools101):

thank u!

hero (hero):

log(x) + log(x+9) = 1 log(x(x+9)) = 1 10^log(x(x+9)) = 10^1 x^2+9x = 10 x^2 + 9x - 10 = 0 x^2+10x-1x-10 = 0 x(x+10)-1(x+10) =0 (x+10)(x-1) = 0 x = -10, x = 1 -10 is invalid, so x = 1 only